First and Second-Order Optimality Conditions for Strong Local Minimum in Control Problems with Pure-State Constraints

Zsolt Páles University of Debrecen, Hungary

Vera Zeidan (speaker) Michigan State University U.S.A.

Abstract

Optimal control problems with pure-state constraints are considered. Unified first and second order conditions for measurable strong local minima are obtained. The first-order condition is the well-known Pontryagin Principle, while the second-order condition is completely new. The method employed is related to the Dubovitskii-Milyutin time transformation with major modifications.